TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-04-07 12:18:00
subject: 3\25 Pt 1 ISS On-Orbit Status 25-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

26 Mar 2003

FOR OFFICIAL USE ONLY

ISS On-Orbit Status 3/25/03

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.

FE-1 Nikolai Budarin had a busy day working on Russian segment (RS)
subsystems. After breakfast, he started out deactivating the Russian
laptop 1 and later turning on laptop 2 with the new software, both on
MCC-M go-ahead.

Then, the Flight Engineer conducted his regular daily checkup of the
BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

A larger task for Budarin was the replacement of a local temperature
sensor switch device (commutator, LKT4B2/TA251) in the Service
Module (SM) with a new spare. There are 11 of these LKTs behind
various SM panels which receive, convert and transmit temperature
sensor data via the local analog telemetry switch (LKA) to the
database. [Before exchanging the TA251 boxes, Budarin pulled their
BITS2-12 onboard telemetry system connections, after the ground had
powered the BITS off. He equipped the new TA251 with the ROM
(read-only memory) unit (TA765B) from the old commutator, then
installed it on its mount. All BITS connectors were then remated for
activation and checkout by the ground.] 

Later, Budarin uninstalled an electronic part (APDG antenna switch)
of the TVS television system's AFU antenna feeder unit in the SM and
replaced it with a new unit.

Another task completed by Nick today was a repair, scheduled for half
an hour, on the PVK-1 "Chibis" suit which is used by cosmonauts as
part of their preparations for returning into gravity. [Using
scissors and gray tape, Nikolai needed to fasten a loose insole with
Velcro to the boot, applying the tape at toe, center and heel. Chibis
is a below-the-waist reduced-pressure device (US: LBNP, lower body
negative pressure), which provides gravity-simulating stress to the
body's cardiovascular/circulatory system for reestablishing the
body's orthostatic tolerance (e.g., the Gauer-Henry reflex) after
long-duration stays in zero-G.] 

For review by the crew, the ground uplinked a list of "yellow-tag"
items, i.e., onboard equipment which should have yellow caution tags
to designate them as equipment for which the safety certification
process for the USOS or RS has not been completed. [This includes
items not certified for ISS operations or cleared for only one
segment but not the other, items which could pose a safety hazard,
and broken or expired items. The yellow-tag list currently has 17
items. MCC-M has now added to the list Progress-delivered bladders
for storing liquid waste or water, which await their certification.] 

FE-2/SO Don Pettit activated and checked out the onboard "Net
Meeting" network for conducting meeting conferences including video
conferencing. 

In Lab and SM, Don Pettit took air flow measurements using the
Velocicalc instrument preparatory to formaldehyde sampling with the
FMK (formaldehyde monitoring kit). [The data required from this test
is the measured flow rate and exact rack location where the flow rate
is equal to or greater than 40 ft/min.]

The crew was also asked by MCC-M to take ppCO2 (carbon dioxide
partial pressure) readings with the U.S. CDMK (CO2 monitoring kit)
near the SM gas analyzer, to assist in troubleshooting the latter.

Pettit continued his work on the MSG (Microgravity Science Glovebox),
today he shutting down the PFMI (pore formation and mobility
investigation) payload. [This involved a few circuit breakers,
experiment power outlets, and video drawer switch throws. After a
number of MSG ground commands, he then was to take the facility to
Standby and power off the MLC (MSG laptop computer).]

In preparation for the upcoming nominal MSG research activity, Pettit
did preparatory work on the InSPACE payload (Investigating the
Structure of Paramagnetic Aggregates from Colloidal Emulsions). [He
inspected five of six coil assemblies, leaving out one damaged
assembly (#CA-002). POC (Payload Operations Center) also requested
some video of the inspections via the Lab camcorder.]

CDR Ken Bowersox unstowed and set up the Foot (foot/ground reaction
forces during space flight) payload equipment, preparing it for EMG
(electromyography, i.e., muscular electric signals recording) and
calibrating it for tomorrow's day-long experiment run.

Bowersox and Pettit completed about two hours of audit work in the
Lab, as first of three parts of a 6-hr. inventory. [Goal of the
time-consuming but very necessary activity is to audit various
stowage locations in "Destiny" and provide the ground information to
update the IMS (inventory management system) database on the basis of
the actual on-orbit situation. The audit is a continuation of the
previous stowage location audits, using an uplinked table organized
the same as for the recent Node audit, with the crew simply filling
in "yes" or "no" for the presence of an item in each location.] 

Don conducted the daily routine maintenance of the SM's SOZh life
support system (including ASU toilet facilities). Nikolai prepared
the daily IMS inventory delta file and completed the regular
inspection of the BRPK-1 air/condensate separator of the SRVK water
processing system. 

All crewmembers completed their daily physical exercise program on
RED (resistive exercise device), TVIS (treadmill with vibration
isolation and stabilization), VELO with load trainer, and on the
freshly repaired CEVIS (cycle ergometer with vibration isolation).

The crew downlinked a TV message for PAO use describing preparations
for and the elements of a crew's return and landing in a Soyuz
vehicle in Kazakhstan. The material will be used as background on
landing day in May. 

Today's targets for the CEO (crew earth observations) program were
Patagonian Glaciers (this was probably the best pass of the day over
southern Patagonia. Looking left of track [northward] for context
views of most of the Southern Patagonian Ice Field. The second pass
later in the day bisected this target region. Crew was to try for
context views of the ice fields and larger glacier by looking either
side of track), Industrialized Southeastern Africa (as cooler high
pressure builds into this region by the time of this pass, aerosol
buildups should have resumed over the industrialized corridor of the
Orange and Vaal River valleys. Looking obliquely left of track to
document the extent and density of smog palls), and High Central
Andean Glaciers (the ISS ascending [south-to-north hemisphere] pass
intersected the crest of the Andes perpendicularly. Crew was to take
advantage of some of the best weather in weeks to try for context
views of the small ice fields here, looking either side of track). 

CEO images can be viewed at the websites
http://eol.jsc.nasa.gov and
http://earthobservatory.nasa.gov

U.S. and Russian Segment Status (as of 1:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS): Elektron O2 generator is powered On (24 Amp). Vozdukh CO2
scrubber is On (manual mode 5). U.S. CDRA CO2 scrubber is Off. TCCS
(trace contaminant control subsystem) is operating. MCA (major
constituents analyzer) is operating. BMP Harmful Impurities unit:
Absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air
conditioner SKV-1 is On; SKV-2 is Off. 

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.